how do you find the complement and supplement of 2pi/21 and 136 degrees ?
1
Expert's answer
2012-04-26T07:52:58-0400
The complement of an acute angle A is the angle 90° – A. So for 2 pi/21 complement is pi/2-2pi/21=(21pi-2pi)/42=19pi/42. Angle 136 degrees is not acute
For any angle A between 0° and 180°, the supplement of A is 180° – A. So for 2pi/21 it's pi-2pi/21=(21pi-2pi)/21=19pi/21
Foe 136 degrees it's 180-136=44 degrees